OpenStudy (anonymous):

How do you simplify (x+3) ! (x+2)!

OpenStudy (anonymous):

They are dividing

OpenStudy (badhi):

is it $$\frac{(x+3)!}{(x+2)!}$$

OpenStudy (anonymous):

yes

OpenStudy (badhi):

If so you can write this as, $$\frac{(x+3)\times(x+2)!}{(x+2)!}$$ Since$$(x+3)!=(x+3)\times (x+2)!$$ Now I guess you can continue

OpenStudy (anonymous):

I dont understand.

OpenStudy (badhi):

$$(x+3)!=(x+3)\times[(x+3)-1]\times[(x+3)-2]\times \cdots \times 1$$ I think you can understand this. $$[(x+3)-1]=(x+2)\\ [(x+3]-2]=[(x+2)-1]\cdots$$ Thus $$(x+3)!=(x+3)\times\underbrace{(x+2)\times[(x+2)-1]\times \cdots \times 1}_{(x+2)!}=(x+3)\times(x+2)!$$ Do you get it now?